kaleb1:

Evaluate the expression below for x = 4, y = –5, and z = 2. 3x^2 -4y+xz A.12 B. 36 C. 52 D. 76

Champion:

Hello @kaleb1 and Welcome! For this we "plug in" the given values into the given equation. For EXAMPLE if x = 4 and they want us to plug it into 3x^2 we would do this: 3(4)^2 = 3*(4*4) = 3*16 = 48 we can do something similar for your question!

Champion:

so we "plug in" x=4 for all the x in your equation y= -5 z=2 and so on it would look something like this: 3*(4)^2 - 4(-5) + (4)(2) then we can simplify (keep in mind the order of operations)

Champion:

This should be able to get you to the answer you need If you're still stuck or need more help, feel free to ask or tag me with @Champion :-)